OpenStudy (anonymous):

Solve the right triangle ABC for the given condition. a=12.1, b=8

OpenStudy (anonymous):

a^2 + b^2 = c^2 12.1^2 + 8^2 = c^2 c = sqrt( 12.1^2 +8^2) c=14.5